Suppose the body in Sample 16.2 is projected at 24.5 m/s, 60° above the horizontal. Find the a. time of flight, b. range, and c. Maximum height.
1
Expert's answer
2020-04-14T09:39:53-0400

Given:

"v_0=24.5\:\rm m/s,"

"\theta=60\degree\!,"

"g=9.81\:\rm m/s^2."


(a) The time of flight

"t=\frac{2v_0\sin\theta}{g}=\frac{2\times 24.5\sin 60\degree}{9.81}=4.33\:\rm s."

(b) The range

"R=\frac{v_0^2\sin2\theta}{g}=\frac{24.5^2\sin 120\degree}{9.81}=53.0\:\rm m."

(c) The maximum height

"H_{\rm max}=\frac{v_0^2\sin^2\theta}{2g}=\frac{24.5^2\sin^2 60\degree}{2\times 9.81}=22.9\:\rm m."